# MUC: Mixture of Uncalibrated Cameras for Robust 3D Human Body Reconstruction

### v1

@@ 4.2 Comparative Analysis with SOTA Methods
Our comparison includes 3D human body reconstruction methods from 2016 to 2023. Most of them are in the category of single view, while a few are in the multi-view category. Meanwhile, many methods adopt the SMPL model, which can only represent human body pose and body shape. A more sophisticated SMPL-X, as adopted by our method, can reconstruct human body pose, hand pose, body shape and facial expression jointly. Specifically, we can treat SMPLer-X [3] as SOTA method for single-view fusion method and Yu et al. [30] as SOTA method for multi-view fusion method. To eliminate differences between using different templates, we not only evaluated using the SMPL-X template but also utilized the conversion tool provided by SMPL-X [24] to transform our prediction results into the SMPL template for evaluation.
@@
  • Our Method’s Edge: Compared to other methods, our approach achieves the best multi-view reconstruction performance, with a PA-MPJPE of 31.3mm, surpassing the previous best multi-view method by 1.7mm. Additionally, it maintains a leading position in both PA-MPJPE and PA-MPVPE metrics in single-view scenarios. After converting results to the SMPL model, our method achieves even better performance, with a PA-MPJPE of 26.5mm, far ahead of any previous methods.
  • Model Bias: Under different model types (SMPL and SMPL-X), our approach demonstrates significant performance improvements in multi-view reconstruction compared to single view. Moreover, it achieves a notable enhancement in joints accuracy compared to previous methods. Since our model utilizes the SMPL-X model for both training and prediction, and the vertex count of the SMPL-X model is roughly twice that of the SMPL model, forcibly converting results to the SMPL model introduces some errors in the overall vertex alignment, thus resulting in poorer performance in the PA-MPVPE metric.
@@
Figure 4: Qualitative comparison with SMPLer-X on RICH dataset. “O” stands for the original image. “S” stands for single-view reconstruction result by SMPLer-X. “F” stands for the fusion result of our method. Zoom in for better view.
4.3 Performance across Multi-view Configurations
To evaluate the effectiveness of our calibration-free multi-view fusion method, we first present a frame-based analysis for a video of three cameras. As illustrated in Figure 5, our multi-view method effectively compensates for the target’s movement in the wild, maintaining a stable PA-MPVPE fluctuation. For example, around frame #130, camera #0 exhibits incorrect pose estimation (highlighted in red bounding-box), leading to a significant spike in PA-MPVPE. Similarly, near frame #480, camera #2 shows errors in estimating arms and legs, resulting in another notable increase in PA-MPVPE for the single-view estimation. For both two cases, the fused result can correctly exhibit human body parts.
@@ 4.3 Performance across Multi-view Configurations
The results, depicted in Table 4, emphasize our method’s ability to leverage additional camera angles effectively. In body reconstruction, we observe a significant reduction in PA-MPJPE from 54.96mm with one camera to 30.50mm with four cameras. The PA-MPVPE also improves, decreasing from 54.58mm to 29.98mm as more cameras are incorporated. Samiliar improvement can also be observed with hand and face. Notably, the transition from one to two cameras marked the most substantial improvement, underlining the value of multi-view geometry. Although the incremental gains lessened with each additional camera, the consistent improvement across all metrics reinforces the efficacy of our approach in utilizing multiple uncalibrated views for 3D human body reconstruction in diverse settings.
@@
1 54.96 54.58 9.34 9.38 9.86
2 33.94 33.00 7.87 7.71 3.70
[…]
4 30.50 29.98 7.27 7.12 3.61
4.4 Impact of Joint and Surface Reweighting
Our qualitative evaluation focuses on the influence of the joint reweighting network on the 3D reconstruction of human body poses. Figure 6 visually demonstrates the effect of JRN’s reweighting across different camera views.
In Figure 6, the size of the red circles superimposed on the model’s joints indicates the importance level that JRN assigns to each joint. Larger circles indicate a higher score, suggesting these joints have a more significant influence on the outcome of the multi-view joint reweighting process. The first row in left, presenting a side angle view, shows only the left side of the body’s joints with minor contributions to the final reconstruction. Conversely, the second row, capturing a frontal view, displays larger weights, indicating a more substantial contribution to the final results. By emphasizing the more reliable joints and down-weighting those that are prone to error, the JRN contributes significantly to the robustness and precision of the reconstructed pose.
Figure 6: Visualization of the predicted scores by JRN across views. Larger red circle stands for larger weight. We only select 21 body joints for clear visualization. “W” stands for the visualization of joint weight. “F” stands for the fusion result of our method. Zoom in for better view.
@@ 4.4 Impact of Joint and Surface Reweighting
Human3.6M RICH
JRN SRN PA-MPJPE PA-MPVPE PA-MPJPE PA-MPVPE
33.03 36.83 34.03 30.83
[…]
Hand Face
JRN SRN PA-MPJPE PA-MPVPE PA-MPVPE
7.92 7.70 3.84
7.89 7.67 3.74
7.22 7.06 3.58
5 Conclusion and Discussion
In this paper, we develop a new method to create 3D mesh of people from images taken by cameras that don’t need special setup and calibration. Our method is scalable to arbitrary number of cameras out of the box. This is a big step forward because it skips the complicated and time-consuming steps usually needed to get cameras ready. Our technique cleverly figures out which parts of the images are reliable and uses this information to build a better 3D reconstruction. It works by looking at how confident it is about the position of different body parts from different camera angles. By doing so, we make sure we get the best possible body pose estimation from every camera, no matter how many we use. Our method is not only easier to use in real-life situations but also gives more accurate results than current state-of-the-art methods.

### v2

@@ Introduction
Our comparison includes 3D human body reconstruction methods from 2016 to 2023. Most of them are in the category of single view, while a few are in the multi-view category. Meanwhile, many methods adopt the SMPL model, which can only represent human body pose and body shape. A more sophisticated SMPL-X, as adopted by our method, can reconstruct human body pose, hand pose, body shape and facial expression jointly. Specifically, we can treat SMPLer-X (Cai et al. 2023) as SOTA method for single-view fusion method and PaFF (Jia et al. 2023) as SOTA method for multi-view fusion method. To eliminate differences between using different templates, we not only evaluated using the SMPL-X template but also utilized the conversion tool provided by SMPL-X (Pavlakos et al. 2019) to transform our prediction results into the SMPL template for evaluation.
@@
  • Our Method’s Edge: Our approach outperforms other methods in multi-view reconstruction, achieving a PA-MPJPE of 27.1mm, which surpasses the previous best multi-view method by 1.1mm. It also performs better than models with larger parameter counts, such as SMPLer-X-large. Additionally, it maintains a leading position in both PA-MPJPE and PA-MPVPE metrics in single-view scenarios.
  • Model Bias: Under different model types (SMPL and SMPL-X), our approach demonstrates significant performance improvements in multi-view reconstruction compared to single view. Moreover, it achieves a notable enhancement in joints accuracy compared to previous methods. Since our model utilizes the SMPL-X model for both training and prediction, and the vertex count of the SMPL-X model is roughly twice that of the SMPL model, forcibly converting results to the SMPL model introduces some errors in the overall vertex alignment. Therefore, we only present results in terms of PA-MPJPE.
@@
Method Multi-view
PA-MPJPE PA-MPVPE
SMPLer-X-base (2023) 42.1 37.1
MUC-base 37.5 33.5
Performance across Multi-view Configurations
To evaluate the effectiveness of our calibration-free multi-view fusion method, we first present a frame-based analysis for a video of three cameras. As illustrated in Figure 5, our multi-view method effectively compensates for the target’s movement in the wild, maintaining a stable PA-MPVPE fluctuation.
@@
The results, shown in Tables 3, underscore our method’s effectiveness in utilizing additional camera angles. For body reconstruction, there is a marked decrease in PA-MPJPE when increasing the camera count from one to four. Similar enhancements are observed in the hand and face reconstructions. Notably, the transition from one to two cameras yields the most significant improvement, highlighting the value of multi-view geometry. While the incremental gains diminish with each additional camera, the consistent improvements across all metrics reinforce our approach’s effectiveness in leveraging multiple uncalibrated views for 3D human body reconstruction in various settings. The same conclusions are also demonstrated on the Human3.6M dataset, with results presented in Table 2 of the supplementary material.
@@
1 52.8 47.7 8.4 8.2 4.1
2 43.7 38.9 7.4 7.2 3.5
3 40.4 35.8 7.0 6.8 3.3
4 38.6 34.5 6.8 6.7 3.2
Impact of Joint and Surface Reweighting
Our qualitative evaluation focuses on the influence of the JRN and SRN on the 3D reconstruction of human body poses and shape. Figure 6 visually demonstrates the effect of reweighting across different camera views.
In Figure 6, the size of the red circles superimposed on the model’s joints indicates the importance level that JRN assigns to each joint. Larger circles indicate a higher score, suggesting that these joints exert a more significant influence during the multi-view joint reweighting process. It is evident that with the aid of the joint distance distribution loss, the JRN assigns lower weights to joints with a higher probability of self-occlusion. This strategy effectively harnesses the more accurate parts of predictions from each single viewpoint for fusion. Similarly, predictions using the SRN feature more precise estimates of body shape.
Figure 6: Visualization of the predicted scores by JRN across different views, accompanied by an ablation study on SRN. Larger red circles represent higher weights. For clarity, only 21 body joints are selected for display. “W” stands for the visualization of joint weight.
Human3.6M RICH
JRN SRN PA-MPJPE PA-MPVPE PA-MPJPE PA-MPVPE
33.8 34.8 42.1 37.1
32.5 34.3 39.4 34.2
31.9 33.4 37.5 33.5
Hand Face
JRN SRN PA-MPJPE PA-MPVPE PA-MPVPE
7.2 7.0 5.0
7.0 6.8 3.5
6.9 6.7 3.3
@@
Conclusion and Discussion
In this paper, we present a new method for creating 3D meshes from images taken by uncalibrated cameras, removing the need for complex setup. Our approach scales easily to any number of cameras and uses distance distribution to estimate self-occlusion, enhanced by our SRN for smooth surface fusion. The technique determines the reliability of image segments to improve 3D reconstruction and assesses the positioning confidence of body parts from different views. Currently limited to static images, future updates could extend to video for more precise pose estimation. This method is not only more practical for everyday use but also surpasses current leading techniques in accuracy.